please help...can x=3 be in the domain for the square root of 2-x?
2x^2/x-3? ..
Explain why or why not for each of the above functions? What are the domains of both functions

Answer by Theo(13342)

when x = 3, sqrt(2-x) = sqrt(-1) which is not a real number, so x = 3 can't be in the domain.

when x = 3, (2x^2) / (x-3) = (2x^2) / 0 which is not a real number, so x = 3 can't be in the domain.

square root of a negative number is not allowed.

division by 0 is not allowed.